# When Are Bias-Free ReLU Networks Effectively Linear Networks?

### v2

@@ Appendix B Useful Lemmas
Grönwall’s Inequality (Gronwall, 1919) is a common tool to obtain error bounds when considering approximate differential equations.
@@
The key implication of Condition 3 we exploit is that the input covariance matrix and the input-output correlation averaged over any half space is equal to those averaged over the entire space.
Lemma 11.
Let set ${\mathbb{S}}^{+}$ be an arbitrary half space divided by a hyperplane with normal vector ${\bm{r}}$, namely ${\mathbb{S}}^{+}=\{{\bm{x}}\in{\mathbb{R}}^{D}|{\bm{r}}^{\top}{\bm{x}}>0\}$. Under the first condition in Condition 3, we have $\forall{\bm{r}}$
@@
Under Condition 3, we have $\forall{\bm{r}}$
@@
Thus the average of ${\bm{x}}{\bm{x}}^{\top}$ over ${\mathbb{S}}^{+}$ is equal to the average in the entire space,
@@
The same holds for ${\bm{x}}y({\bm{x}})$. ∎
@@
where the first equality is the law of total expectation and the second equality uses Equation 20. Because the same holds for all rows, Equation 21a is true.
@@
where the first equality is the law of total expectation and the second equality uses Equation 20. Because the same holds for all elements, Equation 21b is true. ∎
@@ C.1 Learning Dynamics: Early Phase
where the equalities hold under Condition 3 as proven by Lemma 12. We solve the approximate early phase dynamics and prove that the errors introduced by the approximation are bounded in Theorem 15.
Theorem 15.
For time $t<\frac{\tau}{s+\operatorname{Tr}{\bm{\Sigma}}}\ln\frac{1}{w_{\mathrm{init}}}$, the solution to Equation 2 starting from small initialization is exponential growth along one direction with small errors
Proof.
@@
Theorem 15 implies two messages. Firstly, the (leaky) ReLU network has the same time-course solution as its linear counterpart except a scale factor determined by $\alpha$, which is consistent with Theorem 7. Secondly, the (leaky) ReLU and linear networks form rank-one weights with small errors in the early phase. We exploit the rank-one weights to reduce the learning dynamics to Equation 10.
@@ C.2 Learning Dynamics: Late Phase
Theorem 7 relies on Conditions 3 and 5 and arrives at three statements: implementing the same function as in Equation 11, having the same weights as in Equation 12, and retaining rank-one weights as Assumption 5. We prove them one by one.
@@
Comparing Equations 32 and 33, we find that when the weights satisfy Assumption 5, the (leaky) ReLU network implements the same function as the linear network except a scale factor
@@
We compute the second terms in the dynamics under Conditions 3 and 5
@@
Under Conditions 3 and 5, Equation 42 can be simplified
@@
Through the same reasoning as the square loss case, Equations 11 and 12 are proved.
@@ Appendix D Two-Layer Networks Learning Dynamics on Orthogonal Datasets
Suppose the low-rank weights in the two-layer bias-free ReLU network trained on the orthogonal dataset are
${\bm{W}}_{1}=\begin{bmatrix}{\bm{W}}_{1}^{(1)}\\ {\bm{W}}_{1}^{(2)}\end{bmatrix}=\begin{bmatrix}{\bm{r}}_{1}^{(1)}\bar{\bm{x}}_{1}^{\top}\\ {\bm{r}}_{1}^{(2)}\bar{\bm{x}}_{2}^{\top}\end{bmatrix},\quad{\bm{W}}_{2}=\begin{bmatrix}{\bm{W}}_{2}^{(1)}&{\bm{W}}_{2}^{(2)}\end{bmatrix}=\begin{bmatrix}{{\bm{r}}_{1}^{(1)}}^{\top}&{{\bm{r}}_{1}^{(2)}}^{\top},\end{bmatrix}$ (44)
where $\bar{\bm{x}}_{1}={\bm{x}}_{1}/\|{\bm{x}}_{1}\|,\bar{\bm{x}}_{2}={\bm{x}}_{2}/\|{\bm{x}}_{2}\|$, and ${\bm{W}}_{1}^{(1)},{\bm{W}}_{1}^{(2)}$ are rank-one block matrices and align with either one of the two data points.
Since the two data points are orthogonal ${\bm{x}}_{1}^{\top}{\bm{x}}_{2}=0$, we have for $\mu=1,2$
@@
$f({\bm{x}}_{\mu};{\bm{W}})\equiv{\bm{W}}_{2}\sigma({\bm{W}}_{1}{\bm{x}}_{\mu})=\sum_{\nu=1}^{2}{\bm{W}}_{2}^{(\nu)}\sigma\left({\bm{W}}_{1}^{(\nu)}{\bm{x}}_{\mu}\right)={\bm{W}}_{2}^{(\mu)}\sigma\left({\bm{W}}_{1}^{(\mu)}{\bm{x}}_{\mu}\right)={\bm{W}}_{2}^{(\mu)}{\bm{W}}_{1}^{(\mu)}{\bm{x}}_{\mu}$
Learning dynamics with summed square loss for each block of weight matrices are
@@
$\tau\dot{\bm{W}}_{1}^{(\mu)}$ $=\sum_{\nu=1}^{2}\sigma^{\prime}\left({\bm{W}}_{1}^{(\mu)}{\bm{x}}_{\nu}\right)\odot{{\bm{W}}_{2}^{(\mu)}}^{\top}\left(y_{\nu}-{\bm{W}}_{2}\sigma({\bm{W}}_{1}{\bm{x}}_{\nu})\right){\bm{x}}_{\nu}^{\top}$
$=\sigma^{\prime}\left({\bm{W}}_{1}^{(\mu)}{\bm{x}}_{\mu}\right)\odot{{\bm{W}}_{2}^{(\mu)}}^{\top}\left(y_{\mu}-{\bm{W}}_{2}\sigma({\bm{W}}_{1}{\bm{x}}_{\mu})\right){\bm{x}}_{\mu}^{\top}$
$={{\bm{W}}_{2}^{(\mu)}}^{\top}\left(y_{\mu}-{\bm{W}}_{2}^{(\mu)}{\bm{W}}_{1}^{(\mu)}{\bm{x}}_{\mu}\right){\bm{x}}_{\mu}^{\top},$
$\tau\dot{\bm{W}}_{2}^{(\mu)}$ $=\sum_{\nu=1}^{2}\left(y_{\nu}-{\bm{W}}_{2}\sigma({\bm{W}}_{1}{\bm{x}}_{\nu})\right)\sigma\left({\bm{W}}_{1}^{(\mu)}{\bm{x}}_{\nu}\right)^{\top}$
$=\left(y_{\mu}-{\bm{W}}_{2}\sigma({\bm{W}}_{1}{\bm{x}}_{\mu})\right)\sigma\left({\bm{W}}_{1}^{(\mu)}{\bm{x}}_{\mu}\right)^{\top}$
$=\left(y_{\mu}-{\bm{W}}_{2}^{(\mu)}{\bm{W}}_{1}^{(\mu)}{\bm{x}}_{\mu}\right){\bm{x}}_{\mu}^{\top}{{\bm{W}}_{1}^{(\mu)}}^{\top}.$
Thus, the learning dynamics of the weights that align with ${\bm{x}}_{\mu}\,(\mu=1,2)$ are
$\tau\dot{\bm{W}}_{1}^{(\mu)}={{\bm{W}}_{2}^{(\mu)}}^{\top}\left(y{\bm{x}}_{\mu}^{\top}-{\bm{W}}_{2}^{(\mu)}{\bm{W}}_{1}^{(\mu)}{\bm{x}}_{\mu}{\bm{x}}_{\mu}^{\top}\right),$ (46a)
$\tau\dot{\bm{W}}_{2}^{(\mu)}=\left(y{\bm{x}}_{\mu}^{\top}-{\bm{W}}_{2}^{(\mu)}{\bm{W}}_{1}^{(\mu)}{\bm{x}}_{\mu}{\bm{x}}_{\mu}^{\top}\right){{\bm{W}}_{1}^{(\mu)}}^{\top}.$ (46b)
This is the same dynamics as training a linear network (with less hidden neurons) on a single data point $({\bm{x}}_{\mu},y_{\mu})$ as given in Equation 3.

### v3

@@ Appendix B Useful Lemmas
B.1 Grönwall’s Inequality
Grönwall’s Inequality (Grönwall, 1919) is a common tool to obtain error bounds when considering approximate differential equations.
@@ B.1 Grönwall’s Inequality
B.2 Data Statistics
We here prove Lemma 5, which states the input covariance matrix and the input-output correlation averaged over any half space are equal to those averaged over the entire space, under Condition 3.
Proof of Lemma 5.
@@ B.2 Data Statistics
By the definition of the conditional expectation, we have that
@@
Thus, the average of ${\bm{x}}{\bm{x}}^{\top}$ over ${\mathbb{S}}^{+}$ is equal to the average in the entire space,
@@
Because ${\bm{x}}y({\bm{x}})$ is also an even function about ${\bm{x}}$ under Condition 3, the same argument holds
[…]
Note that Equation 23 needs both conditions in Condition 3, that are $p({\bm{x}})=p(-{\bm{x}})$ and $y({\bm{x}})=-y(-{\bm{x}})$. However, Equation 22 only needs the first condition, that is $p({\bm{x}})=p(-{\bm{x}})$. ∎
@@
where the first equality is the law of total expectation and the second equality uses Lemma 5. Because the same holds for all rows, Equation 24a is true.
@@
where the first equality is the law of total expectation and the second equality uses Lemma 5. Because the same holds for all elements, Equation 24b is true. ∎
@@ C.1 Learning Dynamics: Early Phase
where the equalities hold under Condition 3 as proven by Lemma 14. For Lemma 6, we solve the approximate early phase dynamics and show that the errors introduced by the approximation are bounded. We presented Lemma 6 in the main text and now prove it.
Proof of Lemma 6.
@@
Lemma 6 implies two messages. Firstly, the (leaky) ReLU network has the same time-course solution as its linear counterpart except a scale factor determined by $\alpha$, which is consistent with Theorem 8. Secondly, the (leaky) ReLU and linear networks form rank-one weights with small errors in the early phase. We exploit the rank-one weights to reduce the learning dynamics to Equation 12.
@@ C.2 Learning Dynamics: Late Phase
Theorem 8 relies on Conditions 3 and 7 and arrives at three statements: implementing the same function as in Equation 13, having the same weights as in Equation 14, and retaining rank-one weights as Assumption 7. We prove them one by one.
@@
Comparing Equations 34 and 35, we find that when the weights satisfy Assumption 7, the (leaky) ReLU network implements the same function as the linear network except a scale factor
@@
We compute the second terms in the dynamics under Conditions 3 and 7
@@
Under Conditions 3 and 7, Equation 44 can be simplified
@@
Through the same reasoning as the square loss case, Equations 13 and 14 are proven.
@@ C.4 Effect of Regularization
We note a concurrent work (Wang et al., 2025) that considers two-layer bias-free ReLU networks with L2 regularization.
@@ Appendix D Two-Layer Networks Learning Dynamics on Orthogonal Datasets
Figure 9: Two-layer bias-free ReLU networks evolve like two linear networks when trained on a dataset with pairwise orthogonal input, that is $\forall\mu\neq\nu,{\bm{x}}_{\mu}^{\top}{\bm{x}}_{\nu}=0$. The dataset has 20 samples, with 10 samples labeled $+1$ and 10 samples labeled $-1$. The input points with $+1$ label have a larger L2 norm than those with $-1$ label. The black curve is the loss of a two-layer bias-free ReLU network trained on all 20 samples. The blue dashed curve is the loss of a two-layer linear network trained on the ten data points with $+1$ label; and the red dashed curve is that trained on the ten data points with $-1$ label. The loss curve of the ReLU network overlaps with those of the two linear networks. (a) Mean square error loss trajectory. (b) Logistic loss trajectory.
Assumption 17.
A two-layer bias-free ReLU network is trained from small initialization on a dataset with orthonormal input, that is ${\bm{x}}_{\mu}^{\top}{\bm{x}}_{\nu}=0$ if $\mu\neq\nu$, and ${\bm{x}}_{\mu}^{\top}{\bm{x}}_{\mu}=1$. We assume that the weights at time $t_{0}$ during training have the following form:
${\bm{W}}_{1}=\begin{bmatrix}{\bm{W}}_{1}^{\mathrm{A}}\\ {\bm{W}}_{1}^{\mathrm{B}}\end{bmatrix}=\begin{bmatrix}u_{\mathrm{A}}{\bm{r}}_{\mathrm{A}}\bar{\bm{\beta}}_{\mathrm{A}}^{\top}\\ u_{\mathrm{B}}{\bm{r}}_{\mathrm{B}}\bar{\bm{\beta}}_{\mathrm{B}}^{\top}\end{bmatrix},\quad{\bm{W}}_{2}=\begin{bmatrix}{\bm{W}}_{2}^{\mathrm{A}}&{\bm{W}}_{2}^{\mathrm{B}}\end{bmatrix}=\begin{bmatrix}u_{\mathrm{A}}{\bm{r}}_{\mathrm{A}}^{\top}&-u_{\mathrm{B}}{\bm{r}}_{\mathrm{B}}^{\top}\end{bmatrix},$ (46)
where $u>0$, $\|{\bm{r}}_{\mathrm{A}}\|=\|{\bm{r}}_{\mathrm{B}}\|=1$ and all entries in ${\bm{r}}_{\mathrm{A}},{\bm{r}}_{\mathrm{B}}$ are non-negative. The vectors ${\bm{\beta}}_{\mathrm{A}},{\bm{\beta}}_{\mathrm{B}}$ are defined as
$\bar{\bm{\beta}}_{\mathrm{A}}$ $=\frac{\sum_{\mu\in{\mathcal{S}}_{\mathrm{A}}}y_{\mu}{\bm{x}}_{\mu}}{\left\|\sum_{\mu\in{\mathcal{S}}_{\mathrm{A}}}y_{\mu}{\bm{x}}_{\mu}\right\|},\quad{\mathcal{S}}_{\mathrm{A}}=\{\mu|y_{\mu}>0\},$ (47a)
$\bar{\bm{\beta}}_{\mathrm{B}}$ $=-\frac{\sum_{\mu\in{\mathcal{S}}_{\mathrm{B}}}y_{\mu}{\bm{x}}_{\mu}}{\left\|\sum_{\mu\in{\mathcal{S}}_{\mathrm{B}}}y_{\mu}{\bm{x}}_{\mu}\right\|},\quad{\mathcal{S}}_{\mathrm{B}}=\{\mu|y_{\mu}<0\}.$ (47b)
Here the set ${\mathcal{S}}_{\mathrm{A}}$ denotes the indices of the data points with positive target output, and ${\mathcal{S}}_{\mathrm{B}}$ denotes those with negative target output.
Proposition 18.
Under Assumption 17, for all $t\geq t_{0}$, the weights of the two-layer bias-free ReLU network will maintain the form in Equation 46 and its learning dynamics is described by
$\tau\dot{\bm{W}}_{1}^{\mathrm{C}}={{\bm{W}}_{2}^{\mathrm{C}}}^{\top}\left({\bm{\beta}}_{\mathrm{C}}^{\top}-{\bm{W}}_{2}^{\mathrm{C}}{\bm{W}}_{1}^{\mathrm{C}}{\bm{\Sigma}}_{\mathrm{C}}\right),\quad\tau\dot{\bm{W}}_{2}^{\mathrm{C}}=\left({\bm{\beta}}_{\mathrm{C}}^{\top}-{\bm{W}}_{2}^{\mathrm{C}}{\bm{W}}_{1}^{\mathrm{C}}{\bm{\Sigma}}_{\mathrm{C}}\right){{\bm{W}}_{1}^{\mathrm{C}}}^{\top},$ (48)
where $\mathrm{C}$ represents either $\mathrm{A}$ or $\mathrm{B}$, and
${\bm{\beta}}_{\mathrm{C}}=\frac{1}{P}\sum_{\mu\in{\mathcal{S}}_{\mathrm{C}}}y_{\mu}{\bm{x}}_{\mu},\quad{\bm{\Sigma}}_{\mathrm{C}}=\frac{1}{P}\sum_{\mu\in{\mathcal{S}}_{\mathrm{C}}}{\bm{x}}_{\mu}{\bm{x}}_{\mu}^{\top}.$ (49)
The dynamics of ${\bm{W}}_{1}^{\mathrm{C}},{\bm{W}}_{2}^{\mathrm{C}}$ is the same as that of a two-layer linear network trained with data points in ${\mathcal{S}}_{\mathrm{C}}$, as given by Equation 3.
Proof.
Because the weights satisfy Equation 46 and the data points are orthonormal, we have
$\forall\mu\in{\mathcal{S}}_{\mathrm{A}},\quad\sigma\left({\bm{W}}_{1}^{\mathrm{A}}{\bm{x}}_{\mu}\right)={\bm{W}}_{1}^{\mathrm{A}}{\bm{x}}_{\mu},\sigma\left({\bm{W}}_{1}^{\mathrm{B}}{\bm{x}}_{\mu}\right)={\bm{0}};$
$\forall\mu\in{\mathcal{S}}_{\mathrm{B}},\quad\sigma\left({\bm{W}}_{1}^{\mathrm{B}}{\bm{x}}_{\mu}\right)={\bm{W}}_{1}^{\mathrm{B}}{\bm{x}}_{\mu},\sigma\left({\bm{W}}_{1}^{\mathrm{A}}{\bm{x}}_{\mu}\right)={\bm{0}}.$
For data points in the dataset, the network computes
@@
$\forall\mu\in{\mathcal{S}}_{\mathrm{A}},\quad f({\bm{x}}_{\mu};{\bm{W}})\equiv{\bm{W}}_{2}\sigma({\bm{W}}_{1}{\bm{x}}_{\mu})={\bm{W}}_{2}^{\mathrm{A}}\sigma\left({\bm{W}}_{1}^{\mathrm{A}}{\bm{x}}_{\mu}\right)+{\bm{W}}_{2}^{\mathrm{B}}\sigma\left({\bm{W}}_{1}^{\mathrm{B}}{\bm{x}}_{\mu}\right)={\bm{W}}_{2}^{\mathrm{A}}{\bm{W}}_{1}^{\mathrm{A}}{\bm{x}}_{\mu},$
$\forall\mu\in{\mathcal{S}}_{\mathrm{B}},\quad f({\bm{x}}_{\mu};{\bm{W}})\equiv{\bm{W}}_{2}\sigma({\bm{W}}_{1}{\bm{x}}_{\mu})={\bm{W}}_{2}^{\mathrm{A}}\sigma\left({\bm{W}}_{1}^{\mathrm{A}}{\bm{x}}_{\mu}\right)+{\bm{W}}_{2}^{\mathrm{B}}\sigma\left({\bm{W}}_{1}^{\mathrm{B}}{\bm{x}}_{\mu}\right)={\bm{W}}_{2}^{\mathrm{B}}{\bm{W}}_{1}^{\mathrm{B}}{\bm{x}}_{\mu}.$
The learning dynamics of ${\bm{W}}_{1}^{\mathrm{A}},{\bm{W}}_{2}^{\mathrm{A}}$ can be calculated as
$\tau\dot{\bm{W}}_{1}^{\mathrm{A}}$ $=\frac{1}{P}\sum_{\mu=1}^{P}\sigma^{\prime}\left({\bm{W}}_{1}^{\mathrm{A}}{\bm{x}}_{\mu}\right)\odot{{\bm{W}}_{2}^{\mathrm{A}}}^{\top}\left(y_{\mu}-{\bm{W}}_{2}\sigma({\bm{W}}_{1}{\bm{x}}_{\mu})\right){\bm{x}}_{\mu}^{\top}$
$=\frac{1}{P}\sum_{\mu\in{\mathcal{S}}_{\mathrm{A}}}^{P}{{\bm{W}}_{2}^{\mathrm{A}}}^{\top}\left(y_{\mu}-{\bm{W}}_{2}^{\mathrm{A}}{\bm{W}}_{1}^{\mathrm{A}}{\bm{x}}_{\mu}\right){\bm{x}}_{\mu}^{\top}$
$={{\bm{W}}_{2}^{\mathrm{A}}}^{\top}\left({\bm{\beta}}_{\mathrm{A}}^{\top}-{\bm{W}}_{2}^{\mathrm{A}}{\bm{W}}_{1}^{\mathrm{A}}{\bm{\Sigma}}_{\mathrm{A}}\right),$
$\tau\dot{\bm{W}}_{2}^{\mathrm{A}}$ $=\sum_{\mu=1}^{P}\left(y_{\mu}-{\bm{W}}_{2}\sigma({\bm{W}}_{1}{\bm{x}}_{\mu})\right)\sigma\left({\bm{W}}_{1}^{\mathrm{A}}{\bm{x}}_{\mu}\right)^{\top}$
$=\sum_{\mu\in{\mathcal{S}}_{\mathrm{A}}}^{P}\left(y_{\mu}-{\bm{W}}_{2}\sigma({\bm{W}}_{1}{\bm{x}}_{\mu})\right)\sigma\left({\bm{W}}_{1}^{\mathrm{A}}{\bm{x}}_{\mu}\right)^{\top}$
$=\left({\bm{\beta}}_{\mathrm{A}}^{\top}-{\bm{W}}_{2}^{\mathrm{A}}{\bm{W}}_{1}^{\mathrm{A}}{\bm{\Sigma}}_{\mathrm{A}}\right){{\bm{W}}_{1}^{\mathrm{A}}}^{\top}.$
This proves Equation 48. We now prove that the weights will maintain the form in Equation 46 for $t\geq t_{0}$. We first calculate a term that we will need:
@@
$\bar{\bm{\beta}}_{\mathrm{A}}^{\top}{\bm{\Sigma}}_{\mathrm{A}}$ $=\frac{\sum_{\mu\in{\mathcal{S}}_{\mathrm{A}}}y_{\mu}{\bm{x}}_{\mu}^{\top}}{\left\|\sum_{\mu\in{\mathcal{S}}_{\mathrm{A}}}y_{\mu}{\bm{x}}_{\mu}\right\|}\frac{\sum_{\mu^{\prime}\in{\mathcal{S}}_{\mathrm{A}}}{\bm{x}}_{\mu^{\prime}}{\bm{x}}_{\mu^{\prime}}^{\top}}{P}$
$=\frac{\sum_{\mu,\mu^{\prime}\in{\mathcal{S}}_{\mathrm{A}}}y_{\mu}{\bm{x}}_{\mu}^{\top}{\bm{x}}_{\mu^{\prime}}{\bm{x}}_{\mu^{\prime}}^{\top}}{\left\|\sum_{\mu\in{\mathcal{S}}_{\mathrm{A}}}y_{\mu}{\bm{x}}_{\mu}\right\|P}$
$=\frac{\sum_{\mu\in{\mathcal{S}}_{\mathrm{A}}}y_{\mu}{\bm{x}}_{\mu}^{\top}}{\left\|\sum_{\mu\in{\mathcal{S}}_{\mathrm{A}}}y_{\mu}{\bm{x}}_{\mu}\right\|P}$
$=\frac{1}{P}\bar{\bm{\beta}}_{\mathrm{A}}^{\top}.$ (50)
By substituting Equation 46 into Equation 48 and using Appendix D, we obtain
$\tau\dot{\bm{W}}_{1}^{\mathrm{A}}$ $=\tau\frac{d}{dt}u_{\mathrm{A}}{\bm{r}}_{\mathrm{A}}\bar{\bm{\beta}}_{\mathrm{A}}^{\top}=u_{\mathrm{A}}{\bm{r}}_{\mathrm{A}}\left({\bm{\beta}}_{\mathrm{A}}^{\top}-u_{\mathrm{A}}{\bm{r}}_{\mathrm{A}}^{\top}u_{\mathrm{A}}{\bm{r}}_{\mathrm{A}}\bar{\bm{\beta}}_{\mathrm{A}}^{\top}{\bm{\Sigma}}_{\mathrm{A}}\right)=u_{\mathrm{A}}{\bm{r}}_{\mathrm{A}}\bar{\bm{\beta}}_{\mathrm{A}}^{\top}\left(\|{\bm{\beta}}_{\mathrm{A}}\|-\frac{1}{P}u_{\mathrm{A}}^{2}\right),$ (51a)
$\tau\dot{\bm{W}}_{2}^{\mathrm{A}}$ $=\tau\frac{d}{dt}u_{\mathrm{A}}{\bm{r}}_{\mathrm{A}}^{\top}=\left({\bm{\beta}}_{\mathrm{A}}^{\top}-u_{\mathrm{A}}{\bm{r}}_{\mathrm{A}}^{\top}u_{\mathrm{A}}{\bm{r}}_{\mathrm{A}}\bar{\bm{\beta}}_{\mathrm{A}}^{\top}{\bm{\Sigma}}_{\mathrm{A}}\right)u_{\mathrm{A}}\bar{\bm{\beta}}_{\mathrm{A}}{\bm{r}}_{\mathrm{A}}^{\top}=u_{\mathrm{A}}\left(\|{\bm{\beta}}_{\mathrm{A}}\|-\frac{1}{P}u_{\mathrm{A}}^{2}\right){\bm{r}}_{\mathrm{A}}^{\top}.$ (51b)
Equating the coefficients in Equation 51 yields a one-dimensional ordinary differential equation about $u_{\mathrm{A}}$:
$\tau\dot{u}_{\mathrm{A}}=u_{\mathrm{A}}\left(\|{\bm{\beta}}_{\mathrm{A}}\|-\frac{1}{P}u_{\mathrm{A}}^{2}\right).$ (52)
Hence, for the ${\bm{W}}_{1}^{\mathrm{A}},{\bm{W}}_{2}^{\mathrm{A}}$ blocks in Equation 46, only the scalar variable $u_{\mathrm{A}}$ evolves while the vectors ${\bm{r}}_{\mathrm{A}},\bar{\bm{\beta}}_{\mathrm{A}}$ stay unchanged.
With the same approach, we can calculate the learning dynamics of ${\bm{W}}_{1}^{\mathrm{B}},{\bm{W}}_{2}^{\mathrm{B}}$
$\tau\dot{\bm{W}}_{1}^{\mathrm{B}}$ $={{\bm{W}}_{2}^{\mathrm{B}}}^{\top}\left({\bm{\beta}}_{\mathrm{B}}^{\top}-{\bm{W}}_{2}^{\mathrm{B}}{\bm{W}}_{1}^{\mathrm{B}}{\bm{\Sigma}}_{\mathrm{B}}\right),$
$\tau\dot{\bm{W}}_{2}^{\mathrm{B}}$ $=\left({\bm{\beta}}_{\mathrm{B}}^{\top}-{\bm{W}}_{2}^{\mathrm{B}}{\bm{W}}_{1}^{\mathrm{B}}{\bm{\Sigma}}_{\mathrm{B}}\right){{\bm{W}}_{1}^{\mathrm{B}}}^{\top}.$
Substituting in Equation 46, we obtain
$\tau\dot{\bm{W}}_{1}^{\mathrm{B}}$ $=\tau\frac{d}{dt}u_{\mathrm{B}}{\bm{r}}_{\mathrm{B}}\bar{\bm{\beta}}_{\mathrm{B}}^{\top}=-u_{\mathrm{B}}{\bm{r}}_{\mathrm{B}}\left({\bm{\beta}}_{\mathrm{B}}^{\top}+u_{\mathrm{B}}{\bm{r}}_{\mathrm{B}}^{\top}u_{\mathrm{B}}{\bm{r}}_{\mathrm{B}}\bar{\bm{\beta}}_{\mathrm{B}}^{\top}{\bm{\Sigma}}_{\mathrm{B}}\right)=u_{\mathrm{B}}{\bm{r}}_{\mathrm{B}}\bar{\bm{\beta}}_{\mathrm{B}}^{\top}\left(\|{\bm{\beta}}_{\mathrm{B}}\|-\frac{1}{P}u_{\mathrm{B}}^{2}\right),$ (53a)
$\tau\dot{\bm{W}}_{2}^{\mathrm{B}}$ $=-\tau\frac{d}{dt}u_{\mathrm{B}}{\bm{r}}_{\mathrm{B}}^{\top}=\left({\bm{\beta}}_{\mathrm{B}}^{\top}+u_{\mathrm{B}}{\bm{r}}_{\mathrm{B}}^{\top}u_{\mathrm{B}}{\bm{r}}_{\mathrm{B}}\bar{\bm{\beta}}_{\mathrm{B}}^{\top}{\bm{\Sigma}}_{\mathrm{B}}\right)u_{\mathrm{B}}\bar{\bm{\beta}}_{\mathrm{B}}{\bm{r}}_{\mathrm{B}}^{\top}=u_{\mathrm{B}}\left(-\|{\bm{\beta}}_{\mathrm{B}}\|+\frac{1}{P}u_{\mathrm{B}}^{2}\right){\bm{r}}_{\mathrm{B}}^{\top}.$ (53b)
Equating the coefficients in Equation 53 yields a one-dimensional ordinary differential equation about $u_{\mathrm{B}}$:
$\tau\dot{u}_{\mathrm{B}}=u_{\mathrm{B}}\left(\|{\bm{\beta}}_{\mathrm{B}}\|-\frac{1}{P}u_{\mathrm{B}}^{2}\right).$ (54)
Hence, for the ${\bm{W}}_{1}^{\mathrm{B}},{\bm{W}}_{2}^{\mathrm{B}}$ blocks in Equation 46, only the scalar variable $u_{\mathrm{B}}$ evolves while the vectors ${\bm{r}}_{\mathrm{B}},\bar{\bm{\beta}}_{\mathrm{B}}$ stay unchanged. Consequently, if the weights satisfy Equation 46 at time $t_{0}$, Equation 46 will hold for all $t\geq t_{0}$. ∎